# EEG-SCMM: Soft Contrastive Masked Modeling for Cross-Corpus EEG-Based Emotion Recognition

### v1

@@ Introduction
Traditional hard CL treats augmented views generated from the same sample as positive pairs, and those from different samples as negative pairs (Chen et al. 2020; Eldele et al. 2021; Yue et al. 2022). During the computation of the contrastive loss, hard values (1 or 0) are assigned to sample pairs, as illustrated in Fig. 1(b) (Hard CL). However, we argue that this approach fails to account for the “short-term continuity” characteristic inherent in human emotions, leading to inaccurate modeling of inter-sample relationships and hindering the generalizability of the learned embeddings.
@@
$D(x_{i},x_{j})=Norm(Dist(x_{i},x_{j}))\in[0,1],$ (2)
where $Dist(\cdot,\cdot)$ is a metric function used to measure the distance between sample pairs, and $Norm(\cdot)$ denotes min-max normalization. In the experiments, we take the negative of cosine similarity as the metric function. Based on the normalized distance $D(x_{i},x_{j})$, we then define a soft assignment $w{(x_{i},x_{j})}$ for each pair of samples ($x_{i}$, $x_{j}$) using the sigmoid function $\sigma(x)=1/(1+{\rm exp}(-x))$:
$w{(x_{i},x_{j})}=2\alpha\cdot\sigma(-D(x_{i},x_{j})/\tau_{w}),$ (3)
where $\alpha\in[0,1]$ is a boundary parameter that controls the upper bound of soft assignments. $\tau_{w}$ is a sharpness parameter, where smaller values of $\tau_{w}$ result in greater differences in $w(\cdot,\cdot)$ between sample pairs, and vice versa. Figure 4 depicts heat maps of soft assignments $w(\cdot,\cdot)$ under different sharpness parameters $\tau_{w}$.
Figure 4: Soft assignments $w(\cdot,\cdot)$ with different $\tau_{w}$. We use heat maps to visualize the soft assignments with different sharpness parameters $\tau_{w}$. For clarity, we set the upper bound $\alpha=1$ in the figure. Best viewed in color.
@@
$p(z_{i},\widetilde{z}_{i})=\frac{{\rm exp}({\rm sim}(z_{i},\widetilde{z}_{i})/\tau_{c})}{\sum_{z^{\prime}\in\mathcal{Z}\backslash\{z_{i}\}}{\rm exp}({\rm sim}(z_{i},z^{\prime})/\tau_{c})},$ (4)
@@
where ${\rm sim(\cdot,\cdot)}$ refers to the cosine similarity, and $\tau_{c}$ is a temperature parameter used to adjust the scale. Based on $p(z_{i},\widetilde{z}_{i})$, the soft contrastive loss is then defined as:
@@
$\begin{split}\mathcal{L}_{C,i}&=-{\rm log}\thinspace p(z_{i},\widetilde{z}_{i})\\ &-\sum_{\begin{subarray}{c}x^{\prime}\in\Omega\backslash\{x_{i},\widetilde{x}_{i}\}\\ z^{\prime}\in\mathcal{Z}\backslash\{z_{i},\widetilde{z}_{i}\}\end{subarray}}w{(x_{i},x^{\prime})}\cdot{\rm log}\thinspace p(z_{i},z^{\prime}),\end{split}$ (5)
@@
where $\Omega=\mathcal{X}\cup\widetilde{\mathcal{X}}$ represents the union of the data spaces of the original and masked samples. By assigning soft weights to different sample pairs, the model is encouraged to better capture the inherent correlations across different samples. During the training process, the final soft contrastive loss $\mathcal{L}_{C}$ is computed by summing and averaging $\mathcal{L}_{C,i}$ across all samples within a mini-batch. Notably, when $\forall{w(x_{i},x^{\prime})}=0$, the soft contrastive loss reduces to the traditional hard contrastive loss.
Aggregate Reconstruction
@@
To further capture the fine-grained relationships between different samples, we incorporate an aggregator for weighted aggregation and reconstruction. Current approaches for masked EEG modeling typically reconstruct the masked portion based on the unmasked portion of a single masked sample (Lan et al. 2024; Pang et al. 2024), following the learning paradigm of MAE (He et al. 2022). However, this single-sample reconstruction strategy overlooks the interactions between samples, leading to a complex and inadequate reconstruction process.
@@
To overcome this limitation, we introduce an aggregator that improves the traditional single-sample reconstruction process. Specifically, we first calculate the cosine similarity between each pair of projected embeddings $(z_{i},z_{j})$ within a mini-batch, resulting in a similarity matrix $\mathcal{S}$. Based on the pairwise similarities in $\mathcal{S}$, the aggregator then performs weighted aggregation of the embedding $h_{i}$. The weighted aggregation process is defined as:
$h_{i}^{r}=\sum_{z^{\prime}\in\mathcal{Z}\backslash\{z_{i}\}}(\frac{{\rm exp}({\rm sim}(z_{i},z^{\prime})/\tau_{c})}{\sum_{z^{\prime\prime}\in\mathcal{Z}\backslash\{z_{i}\}}{\rm exp}({\rm sim}(z_{i},z^{\prime\prime})/\tau_{c})}\cdot h^{\prime}),$ (6)
@@
where $h^{\prime}\in\mathcal{H}\backslash\{h_{i}\}$ represents the encoded embedding corresponding to the projected embedding $z^{\prime}$, and $\mathcal{H}$ denotes the embedding space of the encoder $E$. This approach allows for a more comprehensive reconstruction by aggregating complementary information and incorporating similar features from different samples during the reconstruction process, while suppressing interference from irrelevant noise samples. Finally, the reconstructed embedding $h_{i}^{r}$ is fed into a lightweight decoder $D$ to obtain the reconstructed sample $x_{i}^{r}$. Following the masked modeling paradigm, we use Mean Square Error (MSE) as the reconstruction loss for model optimization, which is defined as:
$\mathcal{L}_{R,i}=\|x_{i}-x_{i}^{r}\|_{2}^{2}.$ (7)
@@
Similar to the soft contrastive loss $\mathcal{L}_{C}$, the final reconstruction loss $\mathcal{L}_{R}$ is computed by summing and averaging $\mathcal{L}_{R,i}$ across all samples within a mini-batch.
The pre-training process of SCMM
@@
During the pre-training process, SCMM is trained by jointly optimizing $\mathcal{L}_{C}$ and $\mathcal{L}_{R}$. The overall pre-training loss is defined as:
$\mathcal{L}=\lambda_{C}\mathcal{L}_{C}+\lambda_{R}\mathcal{L}_{R},$ (8)
@@
where $\lambda_{C}$ and $\lambda_{R}$ are trade-off hyperparameters, which are adaptively adjusted according to the homoscedastic uncertainty of each loss item. Algorithm 1 details the pre-training process of the proposed SCMM.
@@
Input: Unlabeled pre-training dataset $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$. The number of pre-training epochs.
The pre-training process:
@@
2:      /*All operations are performed within a mini-batch*/
@@
8:     Compute the soft contrastive loss $\mathcal{L}_{C}$ using Eq. (5);
@@
12:     Compute the reconstruction loss $\mathcal{L}_{R}$ using Eq. (7);
13:     Compute the pre-training loss $\mathcal{L}$ using Eq. (8);
@@
Experiments
@@
BiDANN (Li et al. 2018) 49.24 / 10.49 60.46 / 11.17 - -
TANN (Li et al. 2021) 58.41 / 07.16 60.75 / 10.61 - -
PR-PL (Zhou et al. 2023)* 61.01 / 10.55 58.74 / 10.71 - -
E2STN (Zhou et al. 2024) 60.51 / 05.41 61.24 / 15.14 - -
SimCLR (Tang et al. 2021)* 47.27 / 08.44 46.89 / 13.41 44.19 / 09.28 42.03 / 10.05
Mixup (Wickstrøm et al. 2022)* 56.86 / 16.83 55.70 / 16.28 54.55 / 17.95 45.79 / 15.16
MAE (He et al. 2022)* 86.49 / 10.57 83.87 / 08.53 86.02 / 08.96 76.74 / 09.18
JCFA (Liu et al. 2024a)* 67.53 / 12.36 62.40 / 07.54 65.99 / 14.04 52.67 / 05.86
Table 1: Experimental results on SEED and SEED-IV under two cross-corpus conditions: same-class and different-class. ”*” indicates that the results are reproduced by ourselves. A $\rightarrow$ B denotes that A is the pre-training dataset, while B is the dataset for model fine-tuning and testing. Best results are highlighted in bold, while the second-best results are underlined.
Hard $\mathcal{L}_{C}$ Soft $\mathcal{L}_{C}$ $\mathcal{L}_{R}$ Same-Class Different-Class
@@
Table 2: Ablation study on SEED and SEED-IV under two cross-corpus conditions: same-class and different-class.
Datasets
We conduct extensive experiments on three publicly available datasets, SEED (Zheng and Lu 2015), SEED-IV (Zheng et al. 2018), and DEAP (Koelstra et al. 2011), to evaluate the model performance of SCMM. These datasets are diverse in terms of EEG equipment, emotional stimuli, data specifications, labeling approaches and subjects, making them well-suited for assessing the model’s efficacy in cross-corpus EEG-based emotion recognition tasks. In our experiments, we use differential entropy (DE) features as inputs. Detailed descriptions of the datasets and pre-processing procedures are provided in Appendix A.
Implementation Details
In the pre-training stage, we set $r$ to 0.5 and $\mu$ to 0.1 for hybrid masking. We set $\alpha$ to 0.5, $\tau_{w}$ to 0.05, and $\tau_{c}$ to 0.5 for soft CL. The Adam optimizer is utilized with a learning rate of 5e-4 and a weight decay of 3e-4. The pre-training process is conducted over 200 epochs with a batch size of 256. We save the parameters $\theta$ from the final epoch as the pre-trained model. In the fine-tuning stage, we input the embeddings $h_{i}$ generated by the encoder $E$ into an emotion classifier consisting of a 2-layer MLP for final emotion recognition. The classifier is optimized using cross-entropy (CE) loss, and the fine-tuning process is conducted over 50 epochs with a batch size of 128. All experiments are conducted using Python 3.9 with PyTorch 1.13 on an NVIDIA GeForce RTX 3090 GPU. Further implementation details and hyperparameter analysis can be found in Appendix B and Appendix E.
Baselines and Experimental Setup
We compare the proposed SCMM against eight competitive baseline methods, including four conventional deep learning methods: BiDANN (Li et al. 2018), TANN (Li et al. 2021), PR-PL (Zhou et al. 2023), and E2STN (Zhou et al. 2024), as well as four self-supervised learning models: SimCLR (Chen et al. 2020; Tang et al. 2021), Mixup (Zhang et al. 2018; Wickstrøm et al. 2022), MAE (He et al. 2022), and JCFA (Liu et al. 2024a). Notably, E2STN and JCFA are specifically designed for cross-corpus EEG-based emotion recognition. To ensure a fair comparison, we adopt a cross-corpus subject-independent protocol in our experiments, consistent with the setup used by JCFA. We evaluate the model performance using the average accuracy and standard deviation (ACC / STD %) across all subjects in the test set. More details about baseline methods and experimental settings are provided in Appendix C.
Experimental Results
SimCLR (Tang et al. 2021)* 53.12 / 13.12 53.75 / 03.61 51.79 / 04.54
Mixup (Wickstrøm et al. 2022)* 48.75 / 14.37 60.62 / 08.68 60.11 / 07.69
MAE (He et al. 2022)* 83.69 / 10.10 72.19 / 07.24 70.50 / 06.30
JCFA (Liu et al. 2024a)* 64.69 / 12.28 61.59 / 06.26 61.06 / 07.37
Table 3: Experimental results for cross-corpus EEG-based emotion recognition on SEED and DEAP.
Evaluation on the SEED and SEED-IV datasets
To fully validate the model performance of SCMM, we conduct extensive experiments under two cross-corpus conditions: (1) same-class and (2) different-class.
(1) For the same-class cross-corpus validation, we conduct two experiments on the SEED and SEED-IV 3-category datasets: pre-training on SEED-IV and fine-tuning on SEED (SEED-IV3 $\rightarrow$ SEED3), and pre-training on SEED and fine-tuning on SEED-IV (SEED3 $\rightarrow$ SEED-IV3). In both experiments, all samples corresponding to fear emotions in the SEED-IV dataset are excluded. The left two columns in Table 1 present the comparison results, indicating that SCMM achieves SOTA performance in both experiments. Specifically, our model achieves classification accuracies of 91.61% and 87.24% with standard deviations of 7.56% and 8.35% in the SEED-IV3 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV3 experiments, outperforming the second-best method MAE by accuracies of 5.12% and 3.37%, respectively.
(2) For the different-class cross-corpus validation, we conduct another two experiments, denoted as SEED-IV4 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV4. These experiments aim to evaluate the model performance when the pre-training and fine-tuning datasets contain different emotion categories. Experimental results in the right two columns of Table 1 demonstrate that SCMM achieves the best performance in both experiments. Specifically, our model achieves classification accuracies of 91.26% and 80.89% with standard deviations of 7.91% and 8.69% in the SEED-IV4 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV4 experiments, surpassing the second-best method MAE by 5.24% and 4.15% in accuracies, respectively. Note that the traditional CL-based models SimCLR, Mixup, and JCFA exhibit relatively poor performance, primarily due to their use of raw EEG signals as inputs. In addition, SCMM significantly outperforms the traditional masked modeling framework MAE, highlighting the superiority of SCMM. In summary, extensive experimental results on the SEED and SEED-IV datasets confirm that our model exhibits superior performance in multiple cross-corpus EEG-based emotion recognition tasks.
Ablation study
To assess the validity of each module in SCMM, we conduct a comprehensive ablation study on the SEED and SEED-IV datasets. Specifically, we design five different models below. (1) Hard $\mathcal{L}_{C}$ trains the model with only hard contrastive loss. (2) Soft $\mathcal{L}_{C}$ trains the model with only soft contrastive loss. (3) $\mathcal{L}_{R}$ trains the model with only reconstruction loss. (4) Hard $\mathcal{L}_{C}$ and $\mathcal{L}_{R}$ trains the model using both hard contrastive loss and reconstruction loss. (5) Soft $\mathcal{L}_{C}$ and $\mathcal{L}_{R}$ trains the model using both soft contrastive loss and reconstruction loss, representing the full model.
Table 2 presents the results of ablation experiments. Specifically, (1) Hard $\mathcal{L}_{C}$ and (2) Soft $\mathcal{L}_{C}$: the first two models compare the performance of hard CL and soft CL. Experimental results indicate that our well-designed soft contrastive learning effectively improves the classification accuracy. (3) $\mathcal{L}_{R}$: the third model removes the soft contrastive loss and trains the model with only reconstruction loss. The results show that the model performs worst in the absence of the contrastive learning constraint. (4) Hard $\mathcal{L}_{C}$ and $\mathcal{L}_{R}$ and (5) Soft $\mathcal{L}_{C}$ and $\mathcal{L}_{R}$: the last two models compare the performance when combining hard and soft contrastive loss with reconstruction loss, respectively. Experimental results demonstrate that the full model achieves the best performance in all experiments, indicating that SCMM significantly enhances the model performance and stability by combining soft contrastive learning and aggregate reconstruction. This improvement is evident under different cross-corpus conditions, demonstrating the feasibility of extending SCMM to real-life aBCI applications.
Discussion
Method ACC / STD (%)
@@ Discussion
ES 90.64 / 07.97 72.75 / 07.06 71.58 / 05.72 -
Table 4: Comparison of soft contrastive learning in the original data space (OS) and embedding space (ES).
Generalization capability analysis
To further validate the generalization capability of SCMM, we conduct additional experiments on the SEED and DEAP datasets under a different-class cross-corpus scenario, denoted as DEAP $\rightarrow$ SEED3, SEED3 $\rightarrow$ DEAP (Valence) and SEED3 $\rightarrow$ DEAP (Arousal). It is notable that the EEG acquisition equipment, emotional stimuli, data specifications, labeling approaches and subjects are completely different between the two datasets. Table 3 presents the experimental results of SCMM compared to existing methods. Specifically, for the DEAP $\rightarrow$ SEED3 experiment, SCMM achieves an accuracy of 91.70% with a standard deviation of 8.07%, outperforming the second-best method MAE by an accuracy of 8.01%. For the SEED3 $\rightarrow$ DEAP (Valence) and SEED3 $\rightarrow$ DEAP (Arousal) experiments, our model achieves classification accuracies of 73.96% and 72.66% with standard deviations of 6.75% and 5.67%, surpassing the second-best method MAE by 1.77% and 2.16% in accuracies. Experimental results demonstrate that the proposed SCMM maintains excellent performance even when the pre-training and fine-tuning datasets are completely different, highlighting its superior generalization capability.
Model performance with limited fine-tuning data

### v2

@@ 4.2. Soft Contrastive Learning
Traditional hard CL treats the same sample and its augmented views as positive pairs, while treating all other samples as negatives (Chen et al., 2020). During the computation of the contrastive loss, hard values (1 or 0) are assigned to sample pairs, as illustrated in Fig. 1(b) (Hard CL). However, we argue that this approach fails to account for the “short-term continuity” characteristic inherent in human emotions, leading to inaccurate modeling of inter-sample relationships and hindering the generalizability of the learned embeddings.
@@
(2) $D(x_{i},x_{j})=Norm(Dist(x_{i},x_{j}))\in[0,1],$
where $Dist(\cdot,\cdot)$ is a metric function used to measure the distance between sample pairs, and $Norm(\cdot)$ denotes min-max normalization. Based on the normalized distance $D(x_{i},x_{j})$, we then define a soft assignment $w{(x_{i},x_{j})}$ for each pair of samples ($x_{i}$, $x_{j}$) using the sigmoid function $\sigma(x)=1/(1+{\rm exp}(-x))$:
(3) $w{(x_{i},x_{j})}=2\alpha\cdot\sigma(-D(x_{i},x_{j})/\tau_{\rm s}),$
where $\alpha\in[0,1]$ is a boundary parameter that controls the upper bound of soft assignments. $\tau_{\rm s}$ is a sharpness parameter, where smaller values of $\tau_{\rm s}$ result in greater differences in $w(\cdot,\cdot)$ between sample pairs, and vice versa. Figure 4 illustrates the differences between soft assignments $w(\cdot,\cdot)$ with different sharpness $\tau_{\rm s}$.
Figure 4. Heatmaps of soft assignments $w(\cdot,\cdot)$ with different sharpness $\tau_{\rm s}$. Best viewed in color. Zoom in for a better view.
[…]
Soft assignments with different sharpness parameters.
@@
(4) $p(z_{i},\widetilde{z}_{i})=\frac{{\rm exp}({\rm sim}(z_{i},\widetilde{z}_{i})/\tau_{\rm c})}{\sum_{z^{\prime}\in\mathcal{Z}\backslash\{z_{i}\}}{\rm exp}({\rm sim}(z_{i},z^{\prime})/\tau_{\rm c})},$
@@
where ${\rm sim(\cdot,\cdot)}$ refers to the cosine similarity, and $\tau_{\rm c}$ is a temperature parameter used to adjust the scale. Based on $p(z_{i},\widetilde{z}_{i})$, the soft contrastive loss is then defined as:
@@
(5) $\begin{split}\mathcal{L}_{{\rm C},i}&=-{\rm log}\thinspace p(z_{i},\widetilde{z}_{i})\\ &-\sum_{\begin{subarray}{c}{\bm{x}}\in\mathcal{X}\backslash\{x_{i},\widetilde{x}_{i}\}\\ {\bm{z}}\in\mathcal{Z}\backslash\{z_{i},\widetilde{z}_{i}\}\end{subarray}}w{(x_{i},{\bm{x}})}\cdot{\rm log}\thinspace p(z_{i},{\bm{z}}),\end{split}$
@@
where $\mathcal{X}=X\cup\widetilde{X}$ represents the union of the data spaces of the original and masked samples. By assigning soft weights to different sample pairs, the model is encouraged to better capture the inherent correlations across different samples. During the training process, the final soft contrastive loss $\mathcal{L}_{\rm C}$ is computed by summing and averaging $\mathcal{L}_{{\rm C},i}$ across all samples within a mini-batch. Notably, when $\forall{w(x_{i},{\bm{x}})}=0$, the soft contrastive loss reduces to the traditional hard contrastive loss.
4.3. Aggregate Reconstruction
@@
To further capture the fine-grained relationships between different samples, we incorporate an aggregator for weighted aggregation and reconstruction. Current approaches for masked EEG modeling typically reconstruct the masked portion based on the unmasked portion of a single masked sample (Lan et al., 2024; Pang et al., 2024), following the learning paradigm of MAE (He et al., 2022a). However, this single-sample reconstruction strategy overlooks the interactions between samples, leading to a complex and inadequate reconstruction process.
@@
To overcome this limitation, we introduce a similarity-aware aggregator that improves the traditional single-sample reconstruction process. Specifically, we first calculate the cosine similarity between each pair of projected embeddings $(z_{i},z_{j})$ within a mini-batch, resulting in a similarity matrix $\mathcal{S}$. Based on the pairwise similarities in $\mathcal{S}$, the aggregator then performs weighted aggregation of the embedding $h_{i}$, defined as:
(6) $h_{i}^{r}=\sum_{z^{\prime}\in\mathcal{Z}\backslash\{z_{i}\}}(\frac{{\rm exp}({\rm sim}(z_{i},z^{\prime})/\tau_{\rm c})}{\sum_{z^{\prime\prime}\in\mathcal{Z}\backslash\{z_{i}\}}{\rm exp}({\rm sim}(z_{i},z^{\prime\prime})/\tau_{\rm c})}\cdot h^{\prime}),$
@@
where $h^{\prime}\in\mathcal{H}\backslash\{h_{i}\}$ represents the encoded embedding corresponding to the projected embedding $z^{\prime}$, and $\mathcal{H}$ denotes the embedding space of the encoder $E$. This approach allows for a more comprehensive reconstruction by aggregating complementary information and incorporating similar features from different samples during the reconstruction process. Finally, the reconstructed embedding $h_{i}^{r}$ is fed into a lightweight decoder $D$ to obtain the reconstructed sample $x_{i}^{r}$. Following the masked modeling paradigm, we use Mean Squared Error (MSE) as the reconstruction loss for model optimization, which is defined as:
(7) $\mathcal{L}_{{\rm R},i}=\|x_{i}-x_{i}^{r}\|_{2}^{2}.$
@@
Similar to the soft contrastive loss $\mathcal{L}_{\rm C}$, the final reconstruction loss $\mathcal{L}_{\rm R}$ is computed by summing and averaging $\mathcal{L}_{{\rm R},i}$ across all samples within a mini-batch.
[…]
Methods Same-Class Different-Class
Transfer Learning
DANN (Ganin et al., 2016)* 51.91 / 09.27 45.90 / 03.60 - -
BiDANN (Li et al., 2018b) 49.24 / 10.49 60.46 / 11.17 - -
TANN (Li et al., 2021) 58.41 / 07.16 60.75 / 10.61 - -
PR-PL (Zhou et al., 2023)* 61.01 / 10.55 58.74 / 10.71 - -
E2STN (Zhou et al., 2025) 60.51 / 05.41 61.24 / 15.14 - -
Self-Supervised Learning
SimCLR (Chen et al., 2020; Tang et al., 2020)* 47.27 / 08.44 46.89 / 13.41 44.19 / 09.28 42.03 / 10.05
Mixup (Zhang, 2017; Wickstrøm et al., 2022)* 56.86 / 16.83 55.70 / 16.28 54.55 / 17.95 45.79 / 15.16
TS-TCC (Eldele et al., 2021)* 55.38 / 11.65 49.43 / 09.44 52.30 / 12.49 44.57 / 06.08
MAE (He et al., 2022a)* 86.49 / 10.57 83.87 / 08.53 86.02 / 08.96 76.74 / 09.18
JCFA (Liu et al., 2024) 67.53 / 12.36 62.40 / 07.54 65.99 / 14.04 52.67 / 05.86
4.4. The Pre-training Process of SCMM
@@
During the pre-training process, SCMM is trained by jointly optimizing $\mathcal{L}_{\rm C}$ and $\mathcal{L}_{\rm R}$. The overall pre-training loss is defined as:
(8) $\mathcal{L}_{\rm pret}=\lambda_{\rm C}\mathcal{L}_{\rm C}+\lambda_{\rm R}\mathcal{L}_{\rm R},$
@@
where $\lambda_{\rm C}$ and $\lambda_{\rm R}$ are trade-off hyperparameters that are adaptively adjusted according to the homoscedastic uncertainty of each loss item (Kendall et al., 2018). Algorithm 1 details the pre-training process of SCMM.
@@
0:
0: - Unlabeled pre-training EEG emotion dataset $X=\{x_{i}\}_{i=1}^{N}$. The number of pre-training epochs.
0:
@@
2:   // All operations are performed within a mini-batch
@@
8:  Compute the soft contrastive loss $\mathcal{L}_{\rm C}$ using Eq. (5);
@@
12:  Compute the reconstruction loss $\mathcal{L}_{\rm R}$ using Eq. (7);
13:  Compute the pre-training loss $\mathcal{L}_{\rm pret}$ using Eq. (8);
@@
5. Experiments
5.1. Datasets
We conduct extensive experiments on three public datasets, SEED (Zheng and Lu, 2015), SEED-IV (Zheng et al., 2018), and DEAP (Koelstra et al., 2011), to evaluate the model performance of SCMM in cross-corpus EEG-based emotion recognition tasks. These datasets are diverse in terms of EEG equipment, emotional stimuli, data specifications, labeling approaches, and subjects, making them well-suited for assessing the model’s efficacy in cross-corpus scenarios. In the experiments, we use 1-s (SEED and DEAP) and 4-s (SEED-IV) DE features as inputs, respectively. Detailed descriptions of the datasets and pre-processing procedures are provided in Appendix A.
5.2. Implementation Details
In the pre-training stage, we set $r$ to 0.5 and $\mu$ to 0.1 for hybrid masking. We use the negative of cosine similarity as $D(\cdot,\cdot)$, and we set $\alpha$ to 0.5, $\tau_{\rm s}$ to 0.05, and $\tau_{\rm c}$ to 0.5 for soft CL. We use Adam as optimizer with an initial learning rate of 5e-4 and a weight decay of 3e-4. The pre-training process is conducted over 200 epochs with a batch size of 256. We save the model parameters $\theta$ from the final epoch as the pre-trained SCMM. In the fine-tuning stage, we input the encoded embedding $h_{i}$ into an emotion classifier consisting of a 2-layer fully connected network for final emotion recognition. For efficient deployment and testing of the model, the pre-trained SCMM is optimized solely using cross-entropy loss during fine-tuning. The fine-tuning process is conducted over 50 epochs with a batch size of 128. All experiments are conducted using Python 3.9 with PyTorch 1.13 on an NVIDIA GeForce RTX 3090 GPU. We release the source code of SCMM at https://github.com/Kyler-RL/SCMM. Further implementation details can be found in Appendix B.
5.3. Baseline Models and Experimental Settings
[…]
Table 2. Experimental results on SEED and DEAP under the different-class cross-corpus condition.
SimCLR (Chen et al., 2020; Tang et al., 2020)* 53.12 / 13.12 53.75 / 03.61 51.79 / 04.54
Mixup (Zhang, 2017; Wickstrøm et al., 2022)* 48.75 / 14.37 60.62 / 08.68 60.11 / 07.69
TS-TCC (Eldele et al., 2021)* 49.37 / 12.50 56.25 / 03.46 54.13 / 04.45
MAE (He et al., 2022a)* 83.69 / 10.10 72.19 / 07.24 70.50 / 06.30
JCFA (Liu et al., 2024) 64.69 / 12.28 61.59 / 06.26 61.06 / 07.37
Table 3. Ablation study on SEED and SEED-IV under same-class and different-class cross-corpus conditions.
@@ 5.3. Baseline Models and Experimental Settings
w/o $\mathcal{L}_{\rm C}$ 89.68 / 09.32 84.24 / 11.90 89.45 / 09.10 77.24 / 09.14
w/o $\mathcal{L}_{\rm R}$ 90.73 / 08.48 85.07 / 11.05 90.96 / 08.36 78.32 / 07.19
SCMM 91.61 / 07.56 87.24 / 08.35 91.26 / 07.91 80.89 / 08.69
5.4. Results Analysis and Comparison
To fully validate the model performance of SCMM, we conduct extensive experiments under two cross-corpus conditions: (1) same-class and (2) different-class. Appendix E presents the within-dataset validation experimental results of SCMM on SEED and SEED-IV. Full results are provided in Appendix F.
(1) Same-Class. We first conduct two experiments on the SEED and SEED-IV 3-category datasets: pre-training on SEED-IV and fine-tuning on SEED (SEED-IV3 $\rightarrow$ SEED3), and pre-training on SEED and fine-tuning on SEED-IV (SEED3 $\rightarrow$ SEED-IV3). In both experiments, all samples corresponding to fear emotions in the SEED-IV dataset are excluded. The left two columns in Table 1 present the comparison results, indicating that SCMM achieves SOTA performance in both experiments. Specifically, our model achieves classification accuracies of 91.61% and 87.24% with standard deviations of 7.56% and 8.35% in the SEED-IV3 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV3 experiments, outperforming the second-best method MAE by accuracies of 5.12% and 3.37%, respectively. In addition, the proposed SCMM is significantly better than transfer learning methods, highlighting its superiority.
(2) Different-Class. We then conduct two experiments on the SEED and SEED-IV 4-category datasets, denoted as SEED-IV4 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV4. These experiments aim to evaluate the model performance when the pre-training and fine-tuning datasets contain different emotion categories. In the experiments, we exclude transfer learning methods since they are not suitable for scenarios where the training and testing datasets contain different emotion categories. Experimental results in the right two columns of Table 1 demonstrate that SCMM achieves the best performance in both experiments. Specifically, our model achieves classification accuracies of 91.26% and 80.89% with standard deviations of 7.91% and 8.69% in the SEED-IV4 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV4 experiments, surpassing the second-best method MAE by 5.24% and 4.15% in accuracies, respectively. Additionally, the traditional CL-based models SimCLR, Mixup, TS-TCC, and JCFA exhibit relatively poor performance across all experiments due to their use of raw EEG signals as inputs.
To further validate the generalization capability of SCMM, we conduct additional experiments on the SEED and DEAP datasets, denoted as DEAP $\rightarrow$ SEED3, SEED3 $\rightarrow$ DEAP (Valence), and SEED3 $\rightarrow$ DEAP (Arousal). Note that the EEG acquisition equipment, emotional stimuli, data specifications, labeling approaches, and subjects are completely different between the two datasets. Table 2 presents the experimental results of SCMM compared to existing methods. Specifically, for the DEAP $\rightarrow$ SEED3 experiment, SCMM achieves an accuracy of 91.70% with a standard deviation of 8.07%, outperforming the second-best method MAE by an accuracy of 8.01%. For the SEED3 $\rightarrow$ DEAP (Valence) and SEED3 $\rightarrow$ DEAP (Arousal) experiments, SCMM achieves classification accuracies of 73.96% and 72.66% with standard deviations of 6.75% and 5.67%, surpassing the second-best method MAE by 1.77% and 2.16% in accuracy. The results show that our model maintains excellent performance even when the pre-training and fine-tuning datasets are completely different, highlighting its superior generalization capability. Further, the comparative analysis of Table 1 and Table 2 reveals that the model performance of SCMM on the same fine-tuning dataset remains stable when pre-training on different datasets. This suggests that our model effectively captures generalizable emotional EEG representations that are robust to dataset variations.
In summary, extensive experimental results on SEED, SEED-IV, and DEAP confirm that our model exhibits superior performance and stability in cross-corpus EEG-based emotion recognition tasks under both same-class and different-class conditions.
Table 4. Comparison of hard CL and soft CL on SEED and SEED-IV under same-class and different-class cross-corpus conditions.
Methods ACC / STD (%)
@@ 5.4. Results Analysis and Comparison
Table 5. Comparison of soft CL in the embedding space (ES) and original data space (OS) on SEED and SEED-IV under same-class and different-class cross-corpus conditions.
Methods ACC / STD (%)
@@
6. Discussions
6.1. Ablation Study
To assess the validity of each module in SCMM, we conduct a comprehensive ablation study on the SEED and SEED-IV datasets. Table 3 presents the results of ablation experiments. Specifically, we design three different models below. (1) w/o $\mathcal{L}_{\rm C}$: the first configuration removes the soft contrastive loss and trains the model using only the reconstruction loss. The results show that the model performs the worst without the contrastive learning constraint. However, it still outperforms the MAE model based on the single-sample reconstruction paradigm. This suggests that our aggregate reconstruction strategy effectively improves the model performance by capturing fine-grained inter-sample relationships. (2) w/o $\mathcal{L}_{\rm R}$: the second configuration removes the reconstruction loss and trains the model using only the soft contrastive loss. Experimental results show that this configuration outperforms the masked modeling approach, indicating that contrastive learning is more effective for EEG-based emotion recognition by capturing high-level semantic features of EEG signals. (3) SCMM: the last configuration trains the model with the soft contrastive loss and reconstruction loss, representing the full SCMM model. The results demonstrate that this configuration achieves the best performance in all experiments, indicating that SCMM significantly enhances the model performance and stability by combining soft contrastive learning and aggregate reconstruction. This improvement is evident under different cross-corpus conditions, demonstrating the feasibility of extending SCMM to real-life aBCI applications.